Please show work

LEE's card store purchase cards from American Greetings for $10 each box. LEE expects to sell 10 boxes of card a week. LEE's store estimates it cost them $5 to place an order, and holding cost is 20 percent of purchase cost per box. It usually takes 5 days to get each order because of its over-demand during the Christmas season.

1. What is the average inventory (nearest value)?

             a. 7.07 boxes.          b. 3.535 boxes          c. 25.495 boxes               d. 50.99 boxes           e. None

2. When should LEE's store place the next order in order not to lose any customer?

             a. when 50 boxes left             b. when 7.14 boxes left       c. when 10 boxes left              d. None

3. What is the annual total cost of carrying the cards?

             a. $14.14                 b. $28.28               c. $50.99               d. $101.981          e. None
Answer #1

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

AVERAGE INVENTORY = EOQ / 2

ROP = DAILY DEMAND * LEAD TIME

ANNUAL CARRYING COST = (EOQ / 2) * HOLDING COST

DEMAND = 10 * 52 WEEKS = 520

CARRYING COST = 10 * 20% = 2

1.EOQ = SQRT(2 * 520 * 5 / 2) = 50.99

AVERAGE INVENTORY = 50.99 / 2 = 25.495

THE CORRECT OPTION IS C.

2. ROP = (10/7) * 5 = 7.14

THE CORRECT OPTION IS C.

3. ANNUAL CARRYING COST = (50.99 / 2) * 2 = 50.99

THE CORRECT OPTION IS C.

DO GIVE A POSITIVE RATING IF YOU THINK THIS ANSWER HELPED